# LinPrim: Linear Primitives for Differentiable Volumetric Rendering

### v1

@@
Abstract.
Volumetric rendering has become central to modern novel view synthesis methods, which use differentiable rendering to optimize 3D scene representations directly from observed views. While many recent works build on NeRF (Mildenhall et al., 2022) or 3D Gaussians (Kerbl et al., 2023), we explore an alternative volumetric scene representation. More specifically, we introduce two new scene representations based on linear primitives—octahedra and tetrahedra—both of which define homogeneous volumes bounded by triangular faces. This formulation aligns naturally with standard mesh-based tools, minimizing overhead for downstream applications. To optimize these primitives, we present a differentiable rasterizer that runs efficiently on GPUs, allowing end-to-end gradient-based optimization while maintaining real-time rendering capabilities. Through experiments on real-world datasets, we demonstrate comparable performance to state-of-the-art volumetric methods while requiring fewer primitives to achieve similar reconstruction fidelity. Our findings provide insights into the geometry of volumetric rendering and suggest that adopting explicit polyhedra can expand the design space of scene representations.
novel view synthesis, 3D reconstruction, scene representations
copyright: nonejournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Rasterizationccs: Computing methodologies Machine learning approachesccs: Computing methodologies Point-based models
Figure 1. We propose LinPrim, a new method that leverages linear primitives - octahedra and tetrahedra - for differentiable volumetric rendering in order to facilitate 3D scene reconstruction for novel view synthesis. To this end, we propose a differentiable rendering pipeline in conjunction with a real-time capable rasterizer tailored to such primitives, thus achieving interactive frame rates for novel-view rendering and showcasing the potential of explicit polyhedral primitives in modern NVS workflows.
1. Introduction
@@
Representations used in NVS approaches often differ from the triangle-mesh-based approaches used in most 3D pipelines. Triangle meshes, while universal in traditional 3D content creation, are challenging to optimize directly for NVS tasks. Consequently, exploring NVS representations that combine desirable properties of high visual fidelity, performance, stability, and integration into downstream applications is crucial. Although it is possible to convert between most common NVS and traditional representations, doing so adds complexity and usually degrades visual quality (Tang et al., 2023; Huang et al., 2024). This limits the practicality of 3D reconstructions derived from images and consequently hinders their broader adoption.
The strongest visual quality in NVS applications is currently achieved by approaches relying on Neural Radiance Fields (NeRF) (Mildenhall et al., 2022; Barron et al., 2023, 2021, 2022). Although the results are impressive, NeRFs encode an implicit function, making them suboptimal for downstream applications; furthermore, the corresponding volumetric rendering process is expensive and often infeasible for real-time settings. More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) introduced a real-time capable, explicit representation that also achieves strong visual performance. The 3DGS representation consists of 3D Gaussian kernels, which are optimized on known views to adjust their position and features. As evidenced by the large body of derivative works relying on 3DGS (Yu et al., 2024; Huang et al., 2024; Mai et al., 2024; Chen et al., 2024b; Tang et al., 2024), the explicit nature of the representation simplifies integration into other pipelines and settings.
Motivated by the scientific curiosity of how simple, bounded primitives can be used to reconstruct high-fidelity real-world scenes, we explore an alternative to 3DGS by employing transparent polyhedra as our scene representation. Specifically, we introduce two novel scene representations based on octahedron and tetrahedron primitives. Each primitive is characterized by a compact set of features that describe its position, vertices, opacity, and view-dependent appearance. These primitives are bounded by triangular faces and of homogeneous density, which makes them well-suited for direct integration into established graphics workflows and provides an intuitive representation that can be easily understood and modified.
@@ 2. Related Works
A major innovation in NVS has been the development of volumetric rendering approaches that represent scenes as continuous radiance fields. The seminal Neural Radiance Fields (Mildenhall et al., 2022) demonstrated how to learn a 5D function, mapping 3D coordinates and viewing directions to color and density, using a neural network. The rendering process used is differentiable, enabling the optimization of the representation given a known set of input views. Since its introduction, many subsequent works have improved on the initial approach.
In particular, Mip-NeRF (Barron et al., 2021, 2022) and Zip-NeRF (Barron et al., 2023) have drastically improved the visual quality by introducing multi-scale representations, anti-aliasing, and learned compression techniques. Instant-NGP (Müller et al., 2022) takes a different approach by accelerating both training and inference times via a multiresolution hash-based encoding strategy, allowing near real-time reconstruction and rendering.
Despite these advancements, NeRF-based methods remain primarily implicit, often resulting in comparatively high rendering and optimization costs. Their implicit nature can limit usability in downstream tasks. Since geometry and appearance are learned as continuous fields, no explicit surface representation is available by default. Using NeRF-based representations in downstream applications thus relies on specialized approaches or expensive post-processing (Wang et al., 2023; Tang et al., 2023).
@@
Point-based rendering has traditionally been seen as an alternative to mesh-based pipelines due to its simplicity in handling geometry and varying levels of surface detail (Zwicker et al., 2001; Gross and Pfister, 2011). More recently, differentiable point-based methods have emerged, enabling end-to-end optimization of both geometry and appearance from image supervision (Xu et al., 2022; Kopanas et al., 2021).
The seminal work of 3D Gaussian Splatting (Kerbl et al., 2023) replaces traditional point primitives with elliptical Gaussian kernels. By optimizing the position and shape of the kernels from known views, they can create an explicit 3D reconstruction from known views. Additionally, each Gaussian can be efficiently ”splatted” onto the screen space, allowing for real-time rendering performance.
Building on the foundation of 3D Gaussian Splatting, Mip-Splatting (Yu et al., 2024) tackles aliasing by controlling the maximum frequency of the splats, effectively reducing flickering and dilation artifacts. This is achieved by filtering and scaling Gaussian primitives to better match the level of detail required at different viewing distances, preventing excessively sharp or overlapping splats.
Other works replace the Gaussian distributions with other primitives to achieve improved performance. Beyond Gaussians (Chen et al., 2024a), replaces Gaussian kernels with linear kernels to achieve sharper and more precise results in high-frequency regions. Compared to our work, however, the primitives remain elliptical and non-homogeneous in density. In contrast, EVER (Mai et al., 2024) uses homogeneous density ellipsoids to replace alpha-blending with an exact rendering process, which avoids popping artifacts. 3D Convex Splatting (Held et al., 2024) employs 3D smooth convexes, completely detaching it from the Gaussian context. Their primitives offer improved flexibility and reconstruction quality.
@@
3. Method
3.1. Linear Rendering Primitives
@@
Figure 2. Visualization of the distance features of each primitive type. For octahedra, vertices are always located on axes, with the distance being added in the wrong direction and subtracted in the other. For tetrahedra, no such symmetry exists; instead, each vertex has an individual distance value.
@@
We enforce symmetry by choosing a single distance to describe opposing corners, as demonstrated in Fig. 2. Hence, each pair of opposite vertices shares the same distance from the center, effectively reducing the complexity of shape parameters. Aside from improving stability, this ensures that the center is always the true geometric center of the octahedron, simplifying downstream tasks such as bounding volume calculations and ray space approximation.
@@
Although we focus primarily on octahedra, the same principles can be extended to accommodate many types of homogeneous volumes. As a particular example, we demonstrate the usage of tetrahedron primitives to reconstruct scenes. Contrary to octahedra, even regular tetrahedra are not symmetric along all axes and thus require adjustments to the stored features. In particular, we define four basis vectors, which describe directions from the center on which corners are located. All basis vectors are equally spaced such that they can represent regular tetrahedra. We then once again optimize for the distance between the center and each corner, as visualized in Fig. 2, keeping the process largely the same. Contrary to octahedra, this means that the optimized position can differ from the geometric center of the primitive.
@@
Figure 3. Overview of our method. First, primitives are constructed from SfM points. During the rendering process, primitives are preprocessed to adjust them based on the camera pose. Afterward, a front-to-back sorted list of primitives for each tile is created. Lastly, during rasterization, the list is traversed to determine the pixel colors. By comparing the output to a known view, the primitives can then be adjusted using gradient descent. The gradients are similarly used to adjust the population of primitives dynamically.
3.2. Rendering Process
@@
During preprocessing, we construct each primitive from the aforementioned features (i.e. the position, distances, and rotation). Given the camera pose, we then transform the primitives into the camera’s coordinate system and calculate the view-dependent color based on the viewing direction. Afterwards, we project these primitives into the 3D affine ray space approximation introduced in EWA Splatting (Zwicker et al., 2001, 2002). Because small primitives are needed to achieve high-fidelity reconstructions, and because the approximation error grows with increasing distance from the center points, we found that the resulting error remains negligible while still reducing overall performance overhead. Lastly, we compute each primitive’s projected bounding box in screen space to later derive the pixels from which it is visible.
@@
During rasterization, each pixel ray either intersects a given primitive twice or not at all since the introduced primitives are convex. To find intersections, we use the Möller-Trumbore intersection algorithm (MTIA) (Möller and Trumbore, 2005), which is fully differentiable and thus usable in our gradient-based optimization. If there are no intersections, the primitive is not visible; if there are two, the opacity depends on the distance between the intersection points. As our reconstruction is scale ambiguous, we must normalize the obtained intersection distances to ensure consistent appearance across different scene sizes. We do this by determining the smallest distance $d$ of the primitives and then scaling all distance measurements accordingly in a manner similar to EVER (Mai et al., 2024). I.e. given the opacity of a primitive $\alpha$, we define the density of an octahedron as:
(1) $\sigma(\alpha)=-\frac{\log(1-0.99\cdot\alpha)}{2\cdot\min(d_{x},d_{y},d_{z})}.$
Given the density and the two intersection depths $i_{1},i_{2}$, the opacity of an intersected primitive at a given pixel then follows as
(2) $o(i_{1},i_{2})=1-\exp(-\sigma(\alpha)\cdot(i_{2}-i_{1})).$
@@
Finally, we aggregate the color contributions of all visible primitives via alpha blending in front-to-back order. Because we maintain a sorted list of primitives for each tile, we can stop blending once cumulative opacity reaches a predefined threshold (i.e. $0.999$), improving performance without noticeably affecting quality. While this approach simplifies behavior when primitives overlap, the end result is a high-fidelity rendering that faithfully reproduces fine geometry while being efficient enough for real-time applications.
3.3. Anti-Aliasing
In the context of Gaussian kernels, Mip-Splatting (Yu et al., 2024) introduced two antialiasing methods for Gaussian primitives: a 3D smoothing filter and a 2D Mip filter. In conjunction, they are able to control the maximum frequency of primitives, mitigating aliasing and dilation artifacts.
We adopt their 3D smoothing strategy to limit how small primitives can become based on visibility from the training views. As a result, each primitive is visible from at least one pixel of a training view. However, applying their 2D Mip filter approach is not possible for our primitives because they rely on modifying the screen space Gaussian distribution. Instead, for each primitive, we identify vertices that lie on the bounding box used for tiling and shift them outward, parallel to the screen. This effectively enlarges the bounding box footprint and increases the minimum size at which the primitive appears. Moreover, because the fall-off within each primitive is still treated as linear, expanding the distance from the center raises the opacity along the primitive as it is evaluated over a larger extent. Conceptually, this approach acts as a coarse approximation that serves to band-limit the high frequencies that would otherwise cause visible aliasing. Although this approach is not fully mathematically equivalent to a 2D Mip filter, it captures its function and empirically mirrors its behavior.
3.4. Optimization
We employ differentiable rendering to optimize our primitives’ features from known views. For this, we use the loss formulation of 3DGS (Kerbl et al., 2023), consisting of L1 and SSIM terms:
(3) $\mathcal{L}=0.8\mathcal{L}_{1}+0.2\mathcal{L}_{\text{D-SSIM}}.$
The gradients are then optimized using ADAM (Kingma and Ba, 2017) and backpropagated through the entire rendering pipeline, effectively distributing the updates across the primitive features.
@@ 3.4. Optimization
Figure 4. Visualization of our scene representation. We show all octahedra with an opacity of at least $0.25$ on an optimized Mip-NeRF 360 (2022) Bicycle scene reconstruction.
3.5. Population Control
@@
We initialize a primitive for each Structure-from-Motion (SfM) (Schonberger and Frahm, 2016) point using its position and color. Every primitive is constructed with equal distance features to all corners, using the distance between the corresponding SfM point and its closest neighbor as a starting point. While Gaussian kernels are circular when initialized and thus unaffected by rotations, our kernels are not. Thus, we initialize the primitives with a uniformly random rotation quaternion to create a more homogeneous spatial coverage.
@@
Splitting, on the other hand, usually involves creating two new primitives with smaller sizes and placing them by sampling from the original Gaussian as a PDF. To adjust this for octahedra, since the distances are aligned with the coordinate axes and symmetric, we sample the new positions from a PDF with standard deviations set to the corresponding distance. This way, primitives are more likely to spread along the longer dimensions of the octahedra. For tetrahedra, this is not trivially possible and we instead set the standard deviations to a quarter of the largest distance for all dimensions.
A sample result of the resulting scene representation and population of octahedra can be seen in Figure 4.
4. Experiments
We evaluate our results on Mip-NeRF 360 (Barron et al., 2022) and ScanNet++ v2 (Yeshwanth and Liu et al., 2023). For Mip-NeRF 360, we use all 9 available scenes, while for ScanNet++, we use only the first five NVS scenes with available test views. We train and test on the official splits provided by each dataset and maintain consistent parameters across all evaluated scenes. All approaches are trained for 30k iterations.

### v3

@@
Abstract
Volumetric rendering has become central to modern novel view synthesis methods, which use differentiable rendering to optimize 3D scene representations directly from observed views. While many recent works build on NeRF [18] or 3D Gaussians [13], we explore an alternative volumetric scene representation. More specifically, we introduce two new scene representations based on linear primitives—octahedra and tetrahedra—both of which define homogeneous volumes bounded by triangular faces. To optimize these primitives, we present a differentiable rasterizer that runs efficiently on GPUs, allowing end-to-end gradient-based optimization while maintaining real-time rendering capabilities. Through experiments on real-world datasets, we demonstrate comparable performance to state-of-the-art volumetric methods while requiring fewer primitives to achieve similar reconstruction fidelity. Our findings deepen the understanding of 3D representations by providing insights into the fidelity and performance characteristics of transparent polyhedra and suggest that adopting novel primitives can expand the available design space.  [^1] https://nicolasvonluetzow.github.io/LinPrim/
Figure 1: We present LinPrim, a new take on novel view synthesis, that leverages linear primitives - octahedra and tetrahedra - for differentiable volumetric rendering in order to facilitate 3D scene reconstruction. To this end, we propose a differentiable rendering pipeline in conjunction with a real-time capable rasterizer tailored to such primitives, thus achieving interactive frame rates for novel-view rendering and showcasing the potential of polyhedral primitives in NVS workflows.
1 Introduction
@@
Approaches in NVS utilize a wide range of scene representations that differ significantly from the conventional triangle-mesh-based pipelines used in most 3D applications. While triangle meshes are central to standard 3D content creation, their direct optimization in NVS tasks remains challenging [21, 6]. Instead, it is essential to explore the extensive design space of possible representations to assess how they impact fidelity, performance and stability in NVS settings.
The strongest visual quality in NVS applications is currently achieved by approaches relying on Neural Radiance Fields (NeRF) [18, 3, 1, 2]. Although the results are impressive, NeRFs encode an implicit function, making them suboptimal for downstream applications. For instance, editing or scene animation require elements to be easily identified, accessed, or manipulated, which is inherently difficult with NeRF representations [31, 22]. Additionally, the corresponding volumetric rendering process is computationally expensive, often making real-time applications infeasible. More recently, 3D Gaussian Splatting (3DGS) [13] introduced a real-time capable, explicit representation that also achieves strong visual performance. The 3DGS representation consists of 3D Gaussian kernels, which are optimized on known views to adjust their position and features. As evidenced by the large body of derivative works relying on 3DGS [30, 12, 17, 5, 25], the explicit nature of the representation simplifies integration into other pipelines and settings.
Motivated by the scientific curiosity of how simple, bounded primitives can be used to reconstruct high-fidelity real-world scenes, we explore an alternative to 3DGS by employing transparent polyhedra as our scene representation. Specifically, we introduce two novel scene representations based on octahedron and tetrahedron primitives. Each primitive is characterized by a compact set of features that describe its position, vertices, opacity, and view-dependent appearance. These primitives are bounded by triangular faces and of homogeneous density, which makes them an intuitive representation that can be easily understood and modified.
@@ 2 Related Works
A major innovation in NVS has been the development of volumetric rendering approaches that represent scenes as continuous radiance fields. The seminal Neural Radiance Fields [18] demonstrated how to learn a 5D function, mapping 3D coordinates and viewing directions to color and density, using a neural network. The rendering process used is differentiable, enabling the optimization of the representation given a known set of input views. Since its introduction, many subsequent works have improved on the initial approach. In particular, Mip-NeRF [1, 2] and Zip-NeRF [3] have drastically improved the visual quality by introducing multi-scale representations, anti-aliasing, and learned compression techniques. Instant-NGP [20] takes a different approach by accelerating both training and inference times via a multiresolution hash-based encoding strategy, allowing near real-time reconstruction and rendering.
Despite these advancements, NeRF-based methods remain primarily implicit, often resulting in comparatively high rendering and optimization costs. Their implicit nature can limit usability in downstream tasks. Since geometry and appearance are learned as continuous fields, no explicit surface representation is available by default. Using NeRF-based representations in downstream applications thus relies on specialized approaches or expensive post-processing [27, 26].
@@
Point-based rendering has traditionally been seen as an alternative to mesh-based pipelines due to its simplicity in handling geometry and varying levels of surface detail [32, 10]. More recently, differentiable point-based methods have emerged, enabling end-to-end optimization of both geometry and appearance from image supervision [23, 28, 16].
The seminal work of 3D Gaussian Splatting [13] replaces traditional point primitives with elliptical Gaussian kernels. By optimizing the position and shape of the kernels from known views, they can create an explicit 3D reconstruction from known views. Additionally, each Gaussian can be efficiently "splatted" onto the screen space, allowing for real-time rendering performance. Building on the foundation of 3D Gaussian Splatting, Mip-Splatting [30] tackles aliasing by controlling the maximum frequency of the splats, effectively reducing flickering and dilation artifacts. This is achieved by filtering and scaling Gaussian primitives to better match the level of detail required at different viewing distances, preventing excessively sharp or overlapping splats.
Other works replace the Gaussian distributions with other primitives to achieve improved performance. Beyond Gaussians [4], replaces Gaussian kernels with linear kernels to achieve sharper and more precise results in high-frequency regions. Compared to our work, however, the primitives remain elliptical and non-homogeneous in density. In contrast, EVER [17] uses homogeneous density ellipsoids to replace alpha-blending with an exact rendering process, which avoids popping artifacts. 3D Convex Splatting [11] employs 3D smooth convexes, completely detaching it from the Gaussian context. Their primitives offer improved flexibility and reconstruction quality.
@@
3 Method
3.1 Linear Rendering Primitives
@@
We enforce symmetry by choosing a single distance to describe opposing corners. Hence, each pair of opposite vertices shares the same distance from the center, effectively reducing the complexity of shape parameters. Aside from improving stability, this ensures that the center is always the true geometric center of the octahedron, simplifying downstream tasks such as bounding volume calculations and ray space approximation.
@@
Although we focus primarily on octahedra, the same principles can be extended to accommodate many types of homogeneous volumes. As a particular example, we demonstrate the usage of tetrahedron primitives to reconstruct scenes. Contrary to octahedra, even regular tetrahedra are not symmetric along all axes and thus require adjustments to the stored features. In particular, we define four basis vectors, which describe directions from the center on which corners are located. All basis vectors are equally spaced such that they can represent regular tetrahedra. We then once again optimize for the distance between the center and each corner, keeping the process largely the same. Contrary to octahedra, this means that the optimized position can differ from the geometric center of the primitive.
@@
Figure 2: Overview of our method. Primitives are constructed from SfM points. During rendering, primitives are preprocessed according to the camera pose and sorted front-to-back. Afterwards, the list is traversed and the individual color contributions are blended for each pixel. Through comparison with a known view, the primitives’ features and the population are adjusted based on gradient flow.
3.2 Rendering Process
@@
During preprocessing, we construct each primitive from the aforementioned features (i.e. the position, distances, and rotation). Given the camera pose, we then transform the primitives into the camera’s coordinate system and calculate the view-dependent color based on the viewing direction. Afterwards, we project these primitives into the 3D affine ray space approximation introduced in EWA Splatting [32, 33]. Because small primitives are needed to achieve high-fidelity reconstructions, and because the approximation error grows with increasing distance from the center points, we found that the resulting error remains negligible while still reducing overall performance overhead. Lastly, we compute each primitive’s projected bounding box in screen space to later derive the pixels from which it is visible.
@@
During rasterization, each pixel ray either intersects a given primitive twice or not at all since the introduced primitives are convex. To find intersections, we use the Möller-Trumbore intersection algorithm (MTIA) [19], which is fully differentiable and thus usable in our gradient-based optimization. If there are no intersections, the primitive is not visible; if there are two, the opacity depends on the distance between the intersection points. As our reconstruction is scale ambiguous, we must normalize the obtained intersection distances to ensure consistent appearance across different scene sizes. We do this by determining the smallest distance $d$ of the primitives and then scaling all distance measurements accordingly in a manner similar to EVER [17]. I.e. given the opacity of a primitive $\alpha$, we define the density of an octahedron as:
$\sigma(\alpha)=-\frac{\log(1-0.99\cdot\alpha)}{2\cdot\min(d_{x},d_{y},d_{z})}.$ (1)
@@
Finally, we aggregate the color contributions of all visible primitives via alpha blending in front-to-back order. Because we maintain a sorted list of primitives for each tile, we can stop blending once cumulative opacity reaches a predefined threshold (i.e. $0.999$), improving performance without noticeably affecting quality. While this approach simplifies behavior when primitives overlap, the end result is a high-fidelity rendering that faithfully reproduces fine geometry while being efficient enough for real-time applications, as shown in Appendix A.
3.3 Anti-Aliasing
In the context of Gaussian kernels, Mip-Splatting [30] introduced two antialiasing methods for Gaussian primitives: a 3D smoothing filter and a 2D Mip filter. In conjunction, they are able to control the maximum frequency of primitives, mitigating aliasing and dilation artifacts.
We adopt their 3D smoothing strategy to limit how small primitives can become based on visibility from the training views. As a result, each primitive is visible from at least one pixel of a training view. However, applying their 2D Mip filter approach is not possible for our primitives because they rely on modifying the screen space Gaussian distribution. Instead, for each primitive, we identify vertices that lie on the bounding box used for tiling and shift them outward, parallel to the screen. This effectively enlarges the bounding box footprint and increases the minimum size at which the primitive appears. Moreover, because the fall-off within each primitive is still treated as linear, expanding the distance from the center raises the opacity along the primitive as it is evaluated over a larger extent. Conceptually, this approach acts as a coarse approximation that serves to band-limit the high frequencies that would otherwise cause visible aliasing. Although this approach is not fully mathematically equivalent to a 2D Mip filter, it captures its function and empirically mirrors its behavior. We evaluate the effectiveness of the filters in Appendix B.
3.4 Optimization
We employ differentiable rendering to optimize our primitives’ features from known views. For this, we use the loss formulation of 3DGS [13], consisting of L1 and SSIM terms. The gradients are then optimized using ADAM [15] and backpropagated through the entire rendering pipeline, effectively distributing the updates across the primitive features.
@@ 3.4 Optimization
3.5 Population Control
@@
Figure 3: Visualization of our scene representation. We show all octahedra with an opacity of at least $0.25$ on an optimized Mip-NeRF 360 [2] Bicycle scene reconstruction.
@@
We initialize a primitive for each Structure-from-Motion (SfM) [24] point using its position and color. Every primitive is constructed with equal distance features to all corners, using the distance between the corresponding SfM point and its closest neighbor as a starting point. While Gaussian kernels are circular when initialized and thus unaffected by rotations, our kernels are not. Thus, we initialize the primitives with a uniformly random rotation quaternion to create a more homogeneous spatial coverage.
@@
Splitting, on the other hand, usually involves creating two new primitives with smaller sizes and placing them by sampling from the original Gaussian as a PDF. To adjust this for octahedra, since the distances are aligned with the coordinate axes and symmetric, we sample the new positions from a PDF with standard deviations set to the corresponding distance and apply rotation afterwards. This way, primitives are more likely to spread along the longer dimensions of the octahedra. For tetrahedra, this is not trivially possible and we instead set all standard deviations to half the largest distance. A sample result of the resulting population of octahedra can be seen in Figure 3.
4 Experiments
We evaluate our results on Mip-NeRF 360 [2] and ScanNet++ v2 [29]. For Mip-NeRF 360, we use all 9 available scenes, while for ScanNet++, we use only the first five NVS scenes with available test views. We train and test on the official splits provided by each dataset and maintain consistent parameters across all evaluated scenes. All approaches are trained for 30k iterations.